Suppose there is a monopoly in the market whose consumers are one of two types (A and B). The (inverse) demand of type A consumers are given by P = 60 −QA/4 and the (inverse) demand for type B consumers are given by P = 45 −QB/2 .

Suppose that this monopolist has a constant marginal cost of production of MC(q) = 18.25 + Q/6 . Solve for the equilibrium price charged to each type of consumer and the equilibrium quantity sold to each type of consumer. Calculate total monopolist profits.
Answer #1

Date Date ург. (Р) P = 60-QA AR Р: 45 - 9 = А TR - 60 SA - 9А, Тg = 459р - ? MR = Log bo - SA MK - 45 - во MC = 18 • 25 + 845-BB = 18.25 + OA + COB 45-18-25 = QA & OB + GB 26.75 x 6 = GA +6QB +QB. solving 160.5 = QA + 70B ... equlis). the 2 equatioProfil = TR - TC TR = TRA + TRB = 60 x 59 - 159x59) - 8540 -870-24 = 2669.75 = 45 x 14-5 - (14.5 x 14.5) - 652.5 -105.12 - 5.